# Physics-Informed Geometry-Aware Neural Operator

### v1

@@ Abstract
Engineering design problems often involve solving parametric Partial Differential Equations (PDEs) under variable PDE parameters and domain geometry. Recently, neural operators have shown promise in learning PDE operators and quickly predicting the PDE solutions. However, training these neural operators typically requires large datasets, the acquisition of which can be prohibitively expensive. To overcome this, physics-informed training offers an alternative way of building neural operators, eliminating the high computational costs associated with Finite Element generation of training data. Nevertheless, current physics-informed neural operators struggle with limitations, either in handling varying domain geometries or varying PDE parameters. In this research, we introduce a novel method, the Physics-Informed Geometry-Aware Neural Operator (PI-GANO), designed to simultaneously generalize across both PDE parameters and domain geometries. We adopt a geometry encoder to capture the domain geometry features, and design a novel pipeline to integrate this component within the existing DCON architecture. Numerical results demonstrate the accuracy and efficiency of the proposed method.
@@ 1 Introduction
Numerical solution of partial differential equations (PDEs) has been a very active field of research in the last recent, when in particular Finite Element Method (FEM) was primarily studied. FEM involves discretizing a continuous function space using a discrete mesh and solving high-dimensional a linear system, which can be computationally demanding Dhatt et al. (2012). This computational cost becomes particularly substantial in tasks requiring repetitive solutions. An example of such tasks is engineering design which necessitates solving parameterized PDEs over a wide range of PDE parameters and domain geometries for design evaluation Almasri et al. (2024).
@@
Even though these methods offer the ability of generalizing to various PDE parameters and geometries, they all are data-intensive Wang et al. (2021). To bypass the need for training datasets, which are typically expensive to obtain, physics-informed neural operators, which are inspired by physics-informed neural networks (PINNs) Raissi et al. (2019), have emerged as an effective approach. This approach integrates the governing PDEs directly into the training process, making it possible to develop a neural operator without any training data obtained from high fidelity simulations, such as FEM. Among these methods, Physics-informed DeepONet (PI-DeepONet) Wang et al. (2021) extends the original DeepONet framework by embedding physical laws directly into the loss function with automatic differentiation Van Merriënboer et al. (2018) during training, which can only be applied on a specific mesh. Physics-informed Fourier Neural Operator (PI-FNO) Li et al. (2021) builds upon the original FNO framework, offering generalization across different PDE parameter representations in a square domain while also incorporating physics-informed training methods. Similarly, physics-informed Wavelet Neural Operator (PI-WNO) Navaneeth et al. (2024) is also proposed based on the architecture of WNO, implementing physics-informed training for WNO Tripura and Chakraborty (2022) using stochastic projection. Physics-informed Deep Compositional Operator Network (PI-DCON) Zhong and Meidani (2024) adopts a pooling layer to capture the global features of the variable PDE parameters, which enable the learning the PDE operator on any given irregular domain geometry. However, all these physics-informed methods struggle with generalization to varying geometries. A significant advancement to address this limitation is the Physics-informed PointNet (PI-PointNet) Kashefi and Mukerji (2022), which integrates a physics-informed training algorithm into PointNet. Nonetheless, PI-PointNet lacks the capability to generalize across PDE parameters, which limits its application in tackling complex problems in engineering design which also involves varying boundary conditions, etc.
@@ 2.2 Parameter encoding with physics-informed DeepONet
Figure 1: The architecture of the DeepONet is shown.
@@ 3 Methodology
In this section, we introduce our proposed Physics-informed Geometry-aware Neural Operator (PI-GANO) for PDE operator learning. Our approach is primarily based on the PI-DCON architecture and draws inspiration from PI-PointNet. We will first explain the architecture of DCON in Section 3.1, followed by a detailed discussion of our inspiration and methodological innovations in Section 3.2. In Section 3.3, we will also highlight the main novelties of our method by comparing it with PI-PointNet.
@@ 3.1 Deep Compositional Operator Network
Our model is built upon on the architecture of Deep Compositional Operator Network Zhong and Meidani (2024). Without loss of generality, let us consider a simplistic case where the variability is only in one parameter, here denoted by $g(\bm{x})$. Our goal is then to find an approximation for the following operator,
@@
This shows that while DCON can still predict solution for any collocation points in a varied or unseen domain, it fails to produce unique predictions for distinct domains. This suggests that the model may exhibit poor performance especially when changes in domain geometry is substantial.
@@ 3.2 Proposed model architecture
We seek to develop a model that make a different prediction at the same spatial coordinate when the domain geometry is modified. To encode the domain geometry into the existing architecture of DCON, we introduce PI-GANO, which is inspired by PI-PointNet. Specifically, we represent a given domain geometry, denoted by $\Omega_{i}$, using a number of collocation points, with coordinates $\{\bm{x}^{i}_{j}\}_{j=1}^{s_{i}}$. We compute the high-dimensional features of each collocation point using an MLP mapping $\mathcal{U}_{G}(\cdot,\theta_{G})$. Then, similarly to PI-PointNet, we add a pooling layer to extract the global feature of the geometry in each embedding dimension. Hence, the global features of the $i$-th domain geometry $G^{i}$ is computed by:
$\bm{G}^{i}=\text{AvgPool}(\{\bm{h}_{j}\}_{j=1}^{s_{i}}),$ (17)
where $\{\bm{h}_{j}\}_{j=1}^{s_{i}}=\{\mathcal{U}_{G}(\bm{x}_{j},\theta_{G})\}_{j=1}^{s_{i}}$, and $\theta_{G}$ are trainable parameters of the MLP.
We consider three options for selecting the collocation points: (1) selecting points only on the boundary segments that vary, (2) selecting points on the entire boundary, and (3) selecting points inside the entire domain. As discussed in Section 4.4, selecting points only on the variable boundary segments led to the best performance. Additionally, we select the Average-pooling layer instead of Max-pooling as it slightly enhanced the model performance in our experiments.
Next, we compute the local coordinate feature of the collocation point on which the PDE solution is to be calculated. Specifically, for $\bm{x}^{i}$ in domain geometry $\Omega_{i}$, this is done via a linear layer parameterized by $W_{t}$ and $B_{t}$ and an activation function $\sigma$:
$\bm{h}^{i}=\sigma(W_{t}\bm{x}^{i}+B_{t}).$ (18)
Then we concatenate the geometry global feature $G^{i}$ with each local coordinate feature $\bm{h}^{i}_{j}$ to obtain the collocation point coordinate global embedding of each geometry $\bm{H}^{i}_{j}$ by:
$\bm{H}^{i}=[\bm{h}^{i}\|\bm{G}^{i}].$ (19)
This combined information $\bm{H}^{i}$ will be input to a series of operator layers to produce the PDE solution. Finally, the PDE prediction of our model is given by:
$u_{i}(\bm{x}^{i}_{j})\approx\mathcal{M}_{\theta}(\bm{x}^{i}_{j},g_{i}(\bm{x}),\Omega_{i})=\text{sum}\{\bm{b}\odot...(\bm{b}\odot(W^{2}_{O}\ \sigma(\bm{b}\odot(W^{1}_{O}[\sigma(W_{t}\bm{x}^{i}+B_{t})\|\bm{G}^{i}]+B^{1}_{O}))+B^{2}_{O}))\},$ (20)
@@ 3.3 Derivative approximation
Figure 3 shows the computation graphs of PI-PointNet (top) and our model (bottom) for PDE solution in a two-dimensional domain. Solid arrows indicate the forward computation steps, and dashed arrows show the backward propagation steps for derivative calculation. If the two MLPs in our model are identical, it appears that our model and PI-PointNet would share the exact same computation path for PDE solution prediction. However, in our model, the computation of local (collocation point-based) features is independent from that for global features. This results in different computation paths for the derivatives of the model. For instance, in PI-PointNet, the first order derivative of $u_{j}$ with respect to coordinate $x_{j}$ is formulated by:
@@
Comparing Equation 21 to Equation 22, it is evident that the differentiation process in our proposed architecture involves calculating only the derivatives of MLPs, whereas the PI-PointNet’s differentiation process also involves taking the derivative of the Max-pooling layer. Based on the Universal Approximation Theorem for neural networks Attali and Pagès (1997), an MLP can approximate any continuous function and its derivatives within certain error bounds. This is while automatic differentiation over the Max-pooling layer, negatively impacts the model’s performance. The numerical results presented in Section 4.3 will further show the differences in performance between these two backpropagation paths.
@@ 4 Numerical results
In the numerical experiments, we use the following default settings unless mentioned otherwise. We use the hyperbolic tangent function (Tanh) Lau and Lim (2018) as our activation function to ensure the smoothness in high-order derivatives of the models. For PI-DCON and our model architecture, we use three operator layers of width 512. For the architecture of PI-PointNet and PI-PointNet*, we use three hidden layers of width 512. Adam is the default optimizer with the following default hyper-parameters: $\beta_{1}$ = 0.9 and $\beta_{2}$ = 0.999 Kingma and Ba (2014). Each model is trained on 70% of the sampled PDE parameters and validated on 10% of the sampled PDE parameters. Well-trained models predict the PDE solution for the remaining 20% of the sampled PDE parameters. There are two hyper-parameters used in model training: learning rate and coordinate sampling size ratio in each epoch for computing the PDE residual. We implemented grid search Bergstra and Bengio (2012) to obtain the best set of hyper-parameters and report the corresponding model performance. The possible values of learning rate are 0.001, 0.0005, 0.0002, and 0.0001, and the possible values of coordinate sampling size ratio are 0.3,0.2,0.1,0.05. Model training is performed on an NVIDIA P100 GPU using a batch size of 20. We collect the parameters of PDEs by Monte Carlo Simulation Mooney (1997) of stochastic processes Parzen (1999) and derive the solution of PDEs based on the Finite Element Method using Matlab Matlab (2012).
@@ 4.1 A modified PI-PointNet as a baseline
In its original form, PI-PointNet is not designed for generalization across different PDE parameters. Therefore, to create a baseline with this generalization capability, we have modified it to include a part which encodes PDE parameter information into the model inputs. In this modified architecture, called PI-PointNet*, if the function values of PDE parameters are available, they are concatenated with the coordinates of the collocation points as the model input. If not available, zeros are concatenated instead. This setup ensures that both the PDE parameters and the domain geometry are provided to the model, allowing the PDE solution to depend on both factors. Figure 4 shows the architecture of PI-PointNet* for the scenario where the only variable PDE parameter is the boundary conditions.
Figure 4: The architecture of the modified physics-informed PointNet (PI-PointNet*) is shown.
@@ 4.2.1 A Darcy flow problem
Let us consider a two-dimensional Darcy flow problem in a five-edge polygon. The steady state solution of the system is described by the following equation:
@@
where $p(x,y)$ is the pressure, $k(x,y)$ is the permeability field, $f(x,y)$ is the source term, and $g(x,y)$ is the prescribed pressure on the domain boundaries $\partial\Omega$. Following the same setting as Lu et al. (2022), we consider $k(\bm{x})=1$ and $f(\bm{x})=10$. To create variable domain geometries, we consider an analytical expressions, where the coordinates of the five vertices of the polygon is varied. To do so, we first draw five random coordinates to be the center points for the five circles of radius of $0.25$ as shown in  5. These five center points are uniformly drawn on the perimeter of the circle, with the radius of $1$ centered at $(0,0)$. Inside each of the five circles, we will randomly generate one point to serve as a vertex of the polygon-shaped domain.
@@
We employ the Gaussian process model of Eq. 24 to generate $N=1000$ different imposed pressure functions. Then, given each of these $N$ sampled functions, $\{g_{i}\}_{i=1}^{N}$, we solve the PDE using the Finite Element Method (FEM) Matlab (2012). Specifically, for the $i$-th realized domain $\Omega_{i}$ and BC function $g_{i}$, let $M_{i}$ be the number of discretization locations on $\partial\Omega_{i}$, and $\bm{X}_{\text{BC},i}$ be the coordinates of these locations, i.e.

### v3

@@ Abstract
Engineering design problems often involve solving parametric Partial Differential Equations (PDEs) under variable PDE parameters and domain geometry. Recently, neural operators have shown promise in learning PDE operators and quickly predicting the PDE solutions. However, training these neural operators typically requires large datasets, the acquisition of which can be prohibitively expensive. To overcome this, physics-informed training offers an alternative way of building neural operators, eliminating the high computational costs associated with Finite Element generation of training data. Nevertheless, current physics-informed neural operators struggle with limitations, either in handling varying domain geometries or varying PDE parameters. In this research, we introduce a novel method, the Physics-Informed Geometry-Aware Neural Operator (PI-GANO), designed to simultaneously generalize across both PDE parameters and domain geometries. We adopt a geometry encoder to capture the domain geometry features, and design a novel pipeline to integrate this component within the existing Deep Compositional Operator Network architecture. Numerical results demonstrate the accuracy and efficiency of the proposed method. All the codes and data related to this work are available on GitHub: https://github.com/WeihengZ/Physics-informed-Neural-Foundation-Operator.
@@ 1 Introduction
The finite element method (FEM) as the predominant high fidelity numerical approach for solving partial differential equations (PDEs) involves discretizing a continuous function space using a discrete mesh and solving a high-dimensional linear system Dhatt et al. (2012). The cost associated with this solution can be prohibitively large particularly in tasks requiring repetitive solutions. An example of such tasks is engineering design which necessitates solving parameterized PDEs over a wide range of PDE parameters and domain geometries for design evaluation Almasri et al. (2024).
@@
Even though these methods offer the ability of generalizing to various PDE parameters and geometries, they all are data-intensive Wang et al. (2021). To bypass the need for training datasets, which are typically expensive to obtain, physics-informed neural operators, which are inspired by physics-informed neural networks (PINNs) Raissi et al. (2019), have emerged as an effective approach. This approach integrates the governing PDEs directly into the training process, making it possible to develop a neural operator without any training data obtained from high fidelity simulations, such as FEM. Among these methods, Physics-informed DeepONet (PI-DeepONet) Wang et al. (2021) extends the original DeepONet framework by embedding physical laws directly into the loss function with automatic differentiation Van Merriënboer et al. (2018) during training, which can only be applied on a specific mesh. Physics-informed Fourier Neural Operator (PI-FNO) Li et al. (2021) builds upon the original FNO framework, offering generalization across different PDE parameter representations in a square domain while also incorporating physics-informed training methods. Similarly, physics-informed Wavelet Neural Operator (PI-WNO) Navaneeth et al. (2024) is also proposed based on the architecture of WNO, implementing physics-informed training for WNO Tripura and Chakraborty (2022) using stochastic projection. Physics-informed Deep Compositional Operator Network (PI-DCON) Zhong and Meidani (2024) adopts a pooling layer to capture the global features of the variable PDE parameters, which enables the learning of the PDE operator on any given irregular domain geometry. However, all these physics-informed methods struggle with generalization to varying geometries. A significant advancement to address this limitation is the Physics-informed PointNet (PI-PointNet) Kashefi and Mukerji (2022), which integrates a physics-informed training algorithm into PointNet. Nonetheless, PI-PointNet lacks the capability to generalize across PDE parameters, which limits its application in tackling complex problems in engineering design which also involves varying boundary conditions, etc.
@@ 2.2 Parameter encoding with physics-informed DeepONet
Figure 1: The architecture of the DeepONet is shown. The input to the branch net is the sampled discrete representations of variable parameters $k(\bm{x})$ and $g(\bm{x})$, with $l$ and $m$ discretization points, respectively.
@@ 3 Methodology
In this section, we introduce our proposed Physics-informed Geometry-aware Neural Operator (PI-GANO) for PDE operator learning. Our approach is primarily based on the PI-DCON architecture and draws inspiration from PI-PointNet. Our main goal in the model architecture design is to enable discretization-independent generalizability for both variable parameters and variable geometry. To do so, we seek to integrate the parameter generalizability of PI-DCON with the geometry generalizability of the PI-PointNet architecture, in a way that facilitates physics-informed training.
In the following sections, we discuss how our architecture is inspired from PI-PointNet and present the mathematical representations of the geometry-coordinate embedding. Then, we discuss the architecture of DCON and its limitation in handling varying geometries in Section 3.2, followed by a detailed discussion of our proposed model architecture in Section 3.3.
3.1 Geometry-coordinate encoding
Inspired by PI-PointNet, our approach aims to integrate the coordinate of collocation points with information on the domain geometry to compute the PDE solution. This enables the model to generalize across different domain geometries. We first present the mathematical representation for the geometry-coordinate embedding that encapsulate both local coordinate features and global geometric features.
We first propose a multi-layer perception layers (MLP) $\mathcal{U}_{G}(\cdot,\theta_{G})$ to capture the feature of the domain geometry. Specifically, we represent a given domain geometry, denoted by $\Omega_{i}$, using a number of collocation points, with coordinates $\{\bm{x}^{i}_{j}\}_{j=1}^{s_{i}}$. We compute the high-dimensional features of each collocation point using the MLP. Then, similarly to PI-PointNet, we use a pooling layer to extract the global feature of the geometry in each embedding dimension. Hence, the global features of the $i$-th domain geometry $\bm{G}^{i}$ is computed by:
$\bm{G}^{i}=\text{AvgPool}(\{\bm{h}_{j}^{i}\}_{j=1}^{s_{i}}),$ (12)
where $\{\bm{h}_{j}^{i}\}_{j=1}^{s_{i}}=\{\mathcal{U}_{G}(\bm{x}_{j}^{i},\theta_{G})\}_{j=1}^{s_{i}}$, and $\theta_{G}$ are trainable parameters of the MLP. We consider three options for selecting the collocation points: (1) selecting points only on the boundary segments that vary, (2) selecting points on the entire boundary, and (3) selecting points inside the entire domain. As discussed in Section 4.4, selecting points only on the variable boundary segments led to the best performance. Additionally, we select the Average-pooling layer instead of other pooling layers as it slightly enhanced the model performance in our experiments. The detailed numerical comparison of the effects of different pooling layers are presented in Section 4.4.
Next, we compute the local coordinate feature of the collocation point on which the PDE solution is to be calculated. Specifically, for $\bm{x}_{j}^{i}$ in domain geometry $\Omega_{i}$, this is done via a linear layer parameterized by $W_{t}$ and $B_{t}$ and an activation function $\sigma$:
$\bm{h}_{j}^{i}=\sigma(W_{t}\bm{x}_{j}^{i}+B_{t}).$ (13)
Then we concatenate the geometry global feature $\bm{G}^{i}$ with each local coordinate feature $\bm{h}_{j}^{i}$ to obtain the geometry-coordinate embedding $\bm{H}^{i}_{j}$ of each collocation point $\bm{x}_{j}^{i}$ in the geometry $\Omega_{i}$ by:
$\bm{H}_{j}^{i}=[\bm{h}_{j}^{i}\|\bm{G}^{i}].$ (14)
Using these geometry-coordinate embeddings $\bm{H}$ and the PDE parameter features, we can employ various novel model architectures to predict the PDE solutions. A schematic of our proposed model is shown in Figure 2. In this work we focus on the case where PDE parameters are not required to be represented using the same number of discretization points, which is especially limiting for variable geometries. Therefore, the numerical experiments are conducted using the PI-DCON architecture, as it is, to the best of the authors’ knowledge, the only physics-informed model capable of generalizing to different function discretizations.
Figure 2: A schematic of how information on the domain geometry can be integrated with the location and parameter embeddings. The variable inputs include $l$ and $m$ discrete values of parameter functions, together with the coordinates of $n$ points on the domain boundary.
@@ 3.2 Deep Compositional Operator Network
Our model is built upon on the architecture of Deep Compositional Operator Network Zhong and Meidani (2024). Without loss of generality, let us consider a simplistic case where the variability is only in boundary conditions, denoted by $g(\bm{x})$. Our goal is then to find an approximation for the following operator,
@@
This shows that while DCON can still predict solution for any collocation points in a varied or unseen domain, it fails to produce unique predictions for distinct domains. This suggests that the model may exhibit poor performance especially when changes in domain geometry are substantial.
@@ 3.3 Proposed model architecture
We seek to develop a model that makes a different prediction at the same spatial coordinate when the domain geometry is modified. To encode the domain geometry into the existing architecture of DCON, we introduce PI-GANO, which combines the representations for geometry-coordinate embedding and DCON.
In our proposed model architecture, the coordinate-geometry embedding $\bm{H}_{j}^{i}$ will be input to a series of operator layers to produce the PDE solution. Finally, the PDE prediction of our model is given by:
$u_{i}(\bm{x}^{i}_{j})\approx\mathcal{M}_{\theta}(\bm{x}^{i}_{j},g_{i}(\bm{x}),\Omega_{i})=\text{sum}\{\bm{b}\odot...(\bm{b}\odot(W^{2}_{O}\ \sigma(\bm{b}\odot(W^{1}_{O}[\sigma(W_{t}\bm{x}_{j}^{i}+B_{t})\|\bm{G}^{i}]+B^{1}_{O}))+B^{2}_{O}))\},$ (20)
@@
Figure 4 shows the computation graphs of PI-PointNet (top) and our model (bottom) for PDE solution in a two-dimensional domain. Solid arrows indicate the forward computation steps, and dashed arrows show the backward propagation steps for derivative calculation. If the two MLPs in our model are identical, it appears that our model and PI-PointNet would share the exact same computation path for PDE solution prediction. However, in our model, the computation of local (collocation point-based) features is independent from that of global features. This results in different computation paths for the derivatives of the model. For instance, in PI-PointNet, the first order derivative of $u_{j}$ with respect to coordinate $x_{j}$ is formulated by:
@@
Comparing Equation 21 to Equation 22, it is evident that the differentiation process in our proposed architecture involves calculating only the derivatives of MLPs, whereas the PI-PointNet’s differentiation process also involves taking the derivative of the Max-pooling layer. Based on the Universal Approximation Theorem for neural networks Attali and Pagès (1997), an MLP can approximate any continuous function and its derivatives within certain error bounds. This is while automatic differentiation over the Max-pooling layer, negatively impacts the model’s performance. The numerical results presented in Section 4.3 will further show the differences in performance between these two back-propagation paths.
@@ 4 Numerical results
In the numerical experiments, we use the following default settings unless mentioned otherwise. We use the hyperbolic tangent function (Tanh) Lau and Lim (2018) as our activation function to ensure the smoothness in high-order derivatives of the models. For PI-DCON and our model architecture, we use three operator layers of width 512. For the architecture of PI-PointNet and PI-PointNet*, we use three hidden layers of width 512. Adam is the default optimizer with the following default hyper-parameters: $\beta_{1}$ = 0.9 and $\beta_{2}$ = 0.999 Kingma and Ba (2014). Each model is trained on 70% of the sampled PDE parameters and validated on 10% of the sampled PDE parameters. Well-trained models predict the PDE solution for the remaining 20% of the sampled PDE parameters. There are two hyper-parameters used in model training: learning rate and coordinate sampling size ratio in each epoch for computing the PDE residual. We implemented grid search Bergstra and Bengio (2012) to obtain the best set of hyper-parameters and report the corresponding model performance. The possible values of learning rate are 0.001, 0.0005, 0.0002, and 0.0001, and the possible values of coordinate sampling size ratio are 0.3,0.2,0.1,0.05. For the remaining hyperparameters related to the architecture of DCON, we selected their values based on the hyperparameter studies presented in Zhong and Meidani (2024). Model training is performed on an NVIDIA P100 GPU using a batch size of 20. We collect the parameters of PDEs by Monte Carlo Simulation Mooney (1997) of stochastic processes Parzen (1999) and derive the solution of PDEs based on the Finite Element Method using Matlab Matlab (2012). For each of the two problems investigated in our work, we generated 500 PDE samples for the experiments. In contrast to the work in Zhong and Meidani (2024), which focuses on mesh generalization, we used a consistent resolution across all FEM results in our study, as our primary focus in this study is on generalization across different domain geometries.
@@ 4.1 A modified PI-PointNet as a baseline
In its original form, PI-PointNet is not designed for generalization across different PDE parameters. Therefore, to create a baseline with this generalization capability, we have modified it to include a part which encodes PDE parameter information into the model inputs. In this modified architecture, called PI-PointNet*, if the function values of PDE parameters are available, they are concatenated with the coordinates of the collocation points as the model input. If not available, zeros are concatenated instead. This setup ensures that both the PDE parameters and the domain geometry are provided to the model, allowing the PDE solution to depend on both factors. Figure 5 shows the architecture of PI-PointNet* for the scenario where the only variable PDE parameter is the boundary conditions, where $n_{1}$ is the number of collocation points on the boundary and $n_{2}$ is the number of collocation points inside the domain.
Figure 5: The architecture of the modified physics-informed PointNet (PI-PointNet*).
@@ 4.2.1 A Darcy flow problem
As the first example, we consider a Darcy flow problem over variable domain geometries and boundary conditions. A motivation for creating a fast geometry-aware solver is to facilitate the engineering design problem that seeks the optimal shape of a filtration system that enhances the fluid flow through the porous media. Let us specifically consider a two-dimensional Darcy flow problem in a five-edge polygon, where the steady state solution is described by,
@@
where $p(x,y)$ is the pressure, $k(x,y)$ is the permeability field, $f(x,y)$ is the source term, and $g(x,y)$ is the prescribed pressure on the domain boundaries $\partial\Omega$. Following the same setting as Lu et al. (2022), we consider $k(\bm{x})=1$ and $f(\bm{x})=10$, where $\bm{x}=(x,y)$ denotes the coordinates. To create variable domain geometries, we consider an analytical expressions, where the coordinates of the five vertices of the polygon is varied. To do so, we first draw five random coordinates to be the center points for the five circles of radius of $0.25$ as shown in  6. These five center points are uniformly drawn on the perimeter of the circle, with the radius of $1$ centered at $(0,0)$. Inside each of the five circles, we will randomly generate one point to serve as a vertex of the polygon-shaped domain.
@@
We employ the Gaussian process model of Eq. 24 to generate $N=500$ different imposed pressure functions. Then, given each of these $N$ sampled functions, $\{g_{i}\}_{i=1}^{N}$, we solve the PDE using the Finite Element Method (FEM) Matlab (2012). Specifically, for the $i$-th realized domain $\Omega_{i}$ and BC function $g_{i}$, let $M_{i}$ be the number of discretization locations on $\partial\Omega_{i}$, and $\bm{X}_{\text{BC},i}$ be the coordinates of these locations, i.e.